please help} solve for m : 4(m-2)=-2(3-m): please help} solve for m : 4(m-2)=-2(3-m)
Answer by stanbon(19743)

solve for m : 4(m-2)=-2(3-m)
----------
4m-8 = -6+2m
2m = 2
m = 1
========
Cheers,
Stan H.